# Overcomplete Tensor Decomposition via Koszul–Young Flattenings

### v1

@@ 3 Proofs for Upper Bounds
[…]
@@
[…]
@@
There is also an algorithmic proof using simultaneous diagonalization, which appears in [JLV23]. Given an arbitrary basis $p^{(1)},\ldots,p^{(r)}$ for $\mathrm{span}\{z^{(\ell)}\otimes z^{(\ell)}\,:\,\ell\in[r]\}$, construct the third-order tensor $R:=\sum_{\ell=1}^{r}e^{(\ell)}\otimes p^{(\ell)}\in\mathbb{R}^{r}\otimes\mathbb{R}^{N}\otimes\mathbb{R}^{N}$ where $e^{(\ell)}$ denotes the $\ell$th standard basis vector. There are coefficients $\beta_{\ell m}$ such that $p^{(\ell)}=\sum_{m=1}^{r}\beta_{\ell m}\cdot z^{(m)}\otimes z^{(m)}$ for each $\ell\in[r]$, where $\beta=(\beta_{\ell m}\,:\,\ell,m\in[r])$ is an invertible matrix. This means $R$ admits the decomposition
@@
Form the $mn\times r$ matrix whose columns are $x^{(\ell)}\otimes y^{(\ell)}$. It suffices to identify a square submatrix using all the columns whose determinant is a nonzero polynomial in the entries of $x^{(\ell)},y^{(\ell)}$. For this, it suffices to demonstrate a choice of the vectors $x^{(\ell)},y^{(\ell)}$ for which $\{x^{(\ell)}\otimes y^{(\ell)}\,:\,\ell\in[r]\}$ are linearly independent. Fix an injection $[r]\to[m]\times[n]$. For each $\ell\in[r]$, if $\ell\mapsto(i,j)$ then set $x^{(\ell)}=e^{(i)}$ and $y^{(\ell)}=e^{(j)}$ where $e^{(i)}$ is the $i$th standard unit basis vector. ∎
@@
This condition is unchanged under the operation that swaps $n_{2}$ with $n_{3}$ and replaces $p$ with $q-p-1$. Therefore, condition (vii) on $M^{\prime}$ also holds under the same condition.
@@
[…]
@@
It suffices to identify a nonzero minor of $L$ that uses all the columns. To this end, we will describe a pairing that assigns every column $(i,j,\ell,U)$ to some row $(m,S)$ in such a way that each row is used 0 or 1 times. Imagine the columns grouped into $\bar{r}q$ “epochs,” one for each $\ell$ value, and each epoch is further subdivided into $(p+1)(q-p-1)$ “pillars,” one for each $(i,j)$ pair. The ordering of the epochs is unimportant, but let’s fix one arbitrary ordering, and similarly for the ordering of pillars within an epoch and the ordering of columns within a pillar. Call the first $\bar{r}(p+1)$ epochs “type 1” and the remaining $\bar{r}(q-p-1)$ epochs “type 2.” We first describe which $S$ value gets assigned to a given column $(i,j,\ell,U)$:
@@
Expand the determinant $\det(L^{\prime})$ as a sum over permutations. Each term in the sum gives a single monomial (multiplied by a coefficient ±$\pm 1$ or 00). Consider the term in the sum corresponding to the pairing of rows and columns described above. We claim that this term produces a nonzero monomial that is unique in the sense that no other term produces a (nonzero) scalar multiple of this monomial. Once we establish this claim, we are done. To prove the claim we need to argue that, given our monomial of interest, it is possible to deduce the unique pairing of rows and columns that produced it. Within a fixed pillar $(i,j,\ell)$, the powers of $a^{(i,j,\ell)}_{k}$ in the monomial reveal which multiset of $k$ values are removed from the $U$’s in this pillar to form the associated $S$’s. By design, there are only 1 or 2 distinct $k$ values, namely $\{i\}$ or $\{i,j\}$ (for a type 1 or type 2 epoch, respectively). In either case, the specific pairing between $U$’s and $S$’s can be deduced (using the fact that a type 2 epoch removes $j$ whenever possible). The $S$’s appearing within a pillar are all distinct (when $j$ is removed, the resulting $S$ contains $i$, but when $i$ is removed, it does not). Now the powers of $b_{m}^{(i,j,\ell)}$ in the monomial reveal which multiset of $m$ values are used in each pillar. Recalling the “first available” rule, this allows the $m$ value for each column to be deduced. This completes the proof that $\det(L^{\prime})$ is a nonzero polynomial.
@@
We first consider the special case where $|S\cap[p+1]|$ takes its maximum possible value, $p-1$. This case needs separate consideration due to the special rule for type 2 epochs: if $|U\cap[p+1]|=p$ then $j$ cannot be removed. Fix $S$ with $|S\cap[p+1]|=p-1$. In a type 1 epoch $\ell$, the number of tuples $(i,j,U)$ assigned to $S$ is $2(q-p-1)$ because there are 2 choices for $i\notin S$, then $q-p-1$ choices for $j$, and then $U=S\sqcup\{i\}$. In a type 2 epoch, we have two different contributions to count. First consider $U$ with $|U\cap[p+1]|=p$, and recall that these are never allowed to remove $j$. This gives $2(q-p-1)$ tuples $(i,j,U)$, since there are again $2$ choices for $i\notin S$, then $q-p-1$ choices for $j$, and $U=S\sqcup\{i\}$. We also need to consider $U$ with $|U\cap[p+1]|=p-1$ (i.e., cases where $j$ is removed). This gives $(p-1)(q-p-2)$ tuples $(i,j,U)$, since there are $p-1$ choices for $i\in S$, then $q-p-2$ choices for $j\notin S$, and $U=S\sqcup\{j\}$ (we need $i\in S$ so that $i\in U$, as required by (34)). In total, the number of times that $S$ is used is
@@
Now consider the remaining case: fix $S$ with $h:=|S\cap[p+1]|<p-1$. In a type 1 epoch $\ell$, the number of tuples $(i,j,U)$ assigned to $S$ is $(p+1-h)(q-p-1)$ because there are $p+1-h$ choices for $i\notin S$, then $q-p-1$ choices for $j$, and $U=S\sqcup\{i\}$. In a type 2 epoch, we again have two contributions to count. First consider $U$ with $|U\cap[p+1]|=h+1$ (i.e., $i$ is removed). This gives $(p+1-h)(q-2p+h-1)$ tuples $(i,j,U)$, since there are $p+1-h$ choices for $i\notin S$, then $q-2p+h-1$ choices for $j\notin S$, and $U=S\sqcup\{i\}$. Now consider $U$ with $|U\cap[p+1]|=h$ (i.e., $j$ is removed). This gives $h(q-2p+h-1)$ tuples $(i,j,U)$, since there are $h$ choices for $i\in S$, then $q-2p+h-1$ choices for $j\notin S$, and $U=S\sqcup\{j\}$. In total, the number of times that $S$ is used is
@@ 4.1 Matrix Decomposition Lemma
Let $s$ be the symbolic rank of $A(a)$, which is also the generic rank of $M(a^{(1)}\otimes b^{(1)}\otimes c^{(1)})$. The entries of $A(a)$ are degree-$1$ homogeneous polynomials in $a$, so by Lemma 4.1 there are vectors of polynomials $f_{1}(a),\ldots,f_{s}$ and $g_{1}(a),\ldots,g_{s}(a)$ such that
@@
where we write, e.g., $H_{0}[f_{t}]$ instead of $H_{0}[f_{t}(a)]$ to emphasize that this is a constant (degree-00) vector. Now we can write
@@
The variables $x$ are naturally partitioned into $X_{1}:=\{a_{i}\,:\,i\in[n_{1}]\}$, $X_{2}:=\{b_{i}\,:\,i\in[n_{2}]\}$, and $X_{3}:=\{c_{i}\,:\,i\in[n_{3}]\}$. For $S\subseteq[3]$, let $H_{S}(\cdot)$ denote the terms that are degree $1$ in $X_{i}$ for each $i\in S$ and degree 00 in $X_{i}$ for each $i\notin S$. Note that all terms in $M(a\otimes b\otimes c)$ are degree $1$ in each of $X_{1},X_{2},X_{3}$, so we can rewrite the above as
@@
Now, for any setting of the variables $x^{(\ell)}$, we bound the rank of each $M_{S}$. First, $\mathrm{rank}(M_{\emptyset})\leq s$ because $\mathrm{colspan}(M_{\emptyset})\subseteq\mathrm{span}\{H_{\emptyset}[f_{t}]\,:\,t\in[s]\}$, noting that $H_{\emptyset}[f_{t}(x)]$ is a constant (degree-00) vector, which we denote $H_{\emptyset}[f_{t}]$. Next, note that $H_{\{1\}}[f_{t}(x)]$ has only “$a$” variables (not “$b$” or “$c$”), which we denote by $H_{\{1\}}[f_{t}(x^{(\ell)})]=H_{\{1\}}[f_{t}(a=a^{(\ell)})]$. Each $H_{\{1\}}[f_{t}(a=a^{(\ell)})]$ lies in $\mathrm{span}\{H_{\{1\}}[f_{t}(a=e^{(j)})]\,:\,j\in[n_{1}]\}$ where $e^{(j)}$ are unit basis vectors. This means $\mathrm{rank}(M_{\{1\}})\leq sn_{1}$ because $\mathrm{colspan}(M_{\{1\}})\subseteq\mathrm{span}\{H_{\{1\}}[f_{t}(a=e^{(j)})]\,:\,t\in[s],\,j\in[n_{1}]\}$. Similarly, $\mathrm{rank}(M_{\{2\}})\leq sn_{2}$ and $\mathrm{rank}(M_{\{3\}})\leq sn_{3}$. Applying the same argument to the row span, we also have $\mathrm{rank}(M_{\{1,2,3\}})\leq s$, $\mathrm{rank}(M_{\{2,3\}})\leq sn_{1}$, $\mathrm{rank}(M_{\{1,3\}})\leq sn_{2}$, and $\mathrm{rank}(M_{\{1,2\}})\leq sn_{3}$. Putting it together,
@@
The proof above is similar to that of Theorem 4.4 of [EGOW18], where they give a bound of $r\leq 8n$ for $n\times n\times n$ tensors. We now explain the source of our improvement from $8n$ to $26n+$ 2. When decomposing $M$ into 8 terms $M_{S}$ for $S\subseteq[3]$, the original proof of [EGOW18] bounds the rank of each term by $sn$. We instead keep track of the specific structure of each term and see that two of them, namely $M_{\emptyset}$ and $M_{\{1,2,3\}}$, admit a better bound of $s$.
@@
The variables $b$ are naturally partitioned into $\bigsqcup_{i\in[k],j\in[d]}B_{ij}$ where $B_{ij}:=\{b_{m}^{(i,j)}\,:\,m\in[n_{i}]\}$. For $S\subseteq[k]\times[d]$, let $H_{S}(\cdot)$ denote the terms that are degree $1$ in $B_{ij}$ for each $(i,j)\in S$ and degree 00 in $B_{ij}$ for each $(i,j)\notin S$. Note that all terms in (35) are degree $1$ in every $B_{ij}$, so we can rewrite the above decomposition as
@@ Appendix A Trivial Flattenings
Since rank (of the $B$ matrix) cannot exceed generic rank (see Section 2.1.2), it suffices to exhibit a single choice of the variables $a^{(i,\ell)}$ such that the $b^{(\ell)}$ vectors are linearly independent. Provided $r\leq\prod_{i\in S}n_{i}$, we can choose the variables $a^{(i,\ell)}$ so that the $b^{(\ell)}$ are distinct standard unit basis vectors. Formally, fix an injection $[r]\to\prod_{i\in S}[n_{i}]$ and write $\ell\mapsto(\phi_{i}(\ell)\,:\,i\in S)$ where $\phi_{i}(\ell)\in[n_{i}]$. Then for $i\in S$, set $a^{(i,\ell)}$ equal to the standard unit basis vector $e^{(\phi_{i}(\ell))}\in\mathbb{R}^{n_{i}}$, so that $b^{(\ell)}=\bigotimes_{i\in S}e^{(\phi_{i}(\ell))}$. Applying the same argument to the $c^{(\ell)}$ vectors, we also incur the requirement $r\leq\prod_{i\in[k]\setminus S}n_{i}$. ∎

### v2

@@ 3 Proofs for Upper Bounds
Using (25), the vectors $z^{(V)}(a^{(\ell)};p,q)\otimes b^{(\ell)}$ for $\ell\in[r]$ lie in $\mathrm{colspan}(M)\cap(Z_{p,q}\otimes\mathbb{R}^{n_{2}})$, so the inclusion $\supseteq$” in Lemma 3.2 holds. For the reverse inclusion, assume on the contrary that $\subseteq$” fails. By assumption (i) we have $a^{(\ell)}_{1}\neq 0$ for all $\ell\in[r]$, so using (25) and Lemma 3.1, the vectors $\{z^{(U)}(a^{(\ell)};p,q)\otimes b^{(\ell)}\,:\,U\subseteq[q],\,|U|=p+1,\,1\in U,\,\ell\in[r]\}$ span $\mathrm{colspan}(M)$; in fact, since the number of these vectors matches $\mathrm{rank}(M)$, they must form a basis for $\mathrm{colspan}(M)$. Define $\mathcal{U}_{1}:=\{U\subseteq[q]\,:\,|U|=p+1,\,1\in U,\,U\neq V\}$. Now, failure of $\subseteq$” is equivalent to the existence of coefficients $\{\gamma_{U,\ell}\in\mathbb{R}\,:\,U\in\mathcal{U}_{1},\,\ell\in[r]\}$, not all zero, such that
@@
For any $z\in\mathcal{X}\cap\mathcal{Y}$, the vector $w:=z\otimes z$ must belong to the subspace $\mathrm{span}\{y\otimes y\,:\,y\in\mathcal{Y}\}$. Since a basis for $\mathcal{Y}$ is given by $y^{(1)},\ldots,y^{(r)}$ with $y^{(\ell)}:=d^{(\ell)}\otimes b^{(\ell)}$ (which are linearly independent by assumption (iii)), a basis for the subspace $\mathrm{span}\{y\otimes y\,:\,y\in\mathcal{Y}\}$ is given by $\{y^{(\ell_{1})}\otimes y^{(\ell_{2})}+y^{(\ell_{2})}\otimes y^{(\ell_{1})}\,:\,1\leq\ell_{1}\leq\ell_{2}\leq r\}$, so we can write
@@
[…]
@@
Form the $mn\times r$ matrix whose columns are $x^{(\ell)}\otimes y^{(\ell)}$. It suffices to identify a square submatrix using all the columns whose determinant is a nonzero polynomial in the entries of $x^{(\ell)},y^{(\ell)}$. For this, it suffices to demonstrate a choice of the vectors $x^{(\ell)},y^{(\ell)}$ for which $\{x^{(\ell)}\otimes y^{(\ell)}\,:\,\ell\in[r]\}$ are linearly independent. Fix an injection $r]\to[m]\times[n]$. For each $\ell\in[r]$, if $\ell\mapsto(i,j)$ then set $x^{(\ell)}=e^{(i)}$ and $y^{(\ell)}=e^{(j)}$ where $e^{(i)}$ is the $i$th standard unit basis vector. ∎
@@
This condition is unchanged under the operation that swaps $n_{2}$ with $n_{3}$ and replaces $p$ with qp1q-p-1. Therefore, condition (vii) on $M^{\prime}$ also holds under the same condition.
@@
It will be convenient to work not with $N$ but with an equivalent matrix, $L$. Recall the derivation of $N$ from Section 3.4. There, we used the columns {U: $\{U\,:\,1\in$ U\} as a basis for the column span of $A(a^{(\ell)})$. By Lemma 3.1, and assuming $a^{(\ell)}_{i}\neq 0$ due to generic components, we can just as easily use the basis of columns $\{U\,:\,i\in U\}$ for any $i\in V:=[p+1]$. To define $L$, we will use a different choice of $i\in[p+1]$ for each $\ell$. The purpose of this is to create symmetry among the elements $i\in[p+1]$, so that element $1$ no longer plays a distinguished role. (However, there remains a distinction between $i\in[p+1]$ versus $i\in[q]\setminus[p+1]$, since we have fixed $V=[p+1]$.)
@@
It suffices to identify a nonzero minor of $L$ that uses all the columns. To this end, we will describe a pairing that assigns every column $(i,j,\ell,U)$ to some row $(m,S)$ in such a way that each row is used 0 or 1 times. Imagine the columns grouped into $\bar{r}q$ “epochs,” one for each $\ell$ value, and each epoch is further subdivided into (p+1)(qp1)(p+1)(q-p-1) “pillars,” one for each $(i,j)$ pair. The ordering of the epochs is unimportant, but let’s fix one arbitrary ordering, and similarly for the ordering of pillars within an epoch and the ordering of columns within a pillar. Call the first $\bar{r}(p+1)$ epochs “type 1” and the remaining $\bar{r}(q-p-1)$ epochs “type 2.” We first describe which $S$ value gets assigned to a given column $(i,j,\ell,U)$:
@@
Expand the determinant $\det(L^{\prime})$ as a sum over permutations. Each term in the sum gives a single monomial (multiplied by a coefficient $\pm 1$ or 0). Consider the term in the sum corresponding to the pairing of rows and columns described above. We claim that this term produces a nonzero monomial that is unique in the sense that no other term produces a (nonzero) scalar multiple of this monomial. Once we establish this claim, we are done. To prove the claim we need to argue that, given our monomial of interest, it is possible to deduce the unique pairing of rows and columns that produced it. Within a fixed pillar $(i,j,\ell)$, the powers of $a^{(i,j,\ell)}_{k}$ in the monomial reveal which multiset of $k$ values are removed from the $U$’s in this pillar to form the associated $S$’s. By design, there are only 1 or 2 distinct $k$ values, namely $\{i\}$ or $\{i,j\}$ (for a type 1 or type 2 epoch, respectively). In either case, the specific pairing between $U$’s and $S$’s can be deduced (using the fact that a type 2 epoch removes $j$ whenever possible). The $S$’s appearing within a pillar are all distinct (when $j$ is removed, the resulting $S$ contains $i$, but when $i$ is removed, it does not). Now the powers of $b_{m}^{(i,j,\ell)}$ in the monomial reveal which multiset of $m$ values are used in each pillar. Recalling the “first available” rule, this allows the $m$ value for each column to be deduced. This completes the proof that $\det(L^{\prime})$ is a nonzero polynomial.
@@
We first consider the special case where $|S\cap[p+1]|$ takes its maximum possible value, p1p-1. This case needs separate consideration due to the special rule for type 2 epochs: if $|U\cap[p+1]|=p$ then $j$ cannot be removed. Fix $S$ with $|S\cap[p+1]|=p-1$. In a type 1 epoch $\ell$, the number of tuples $(i,j,U)$ assigned to $S$ is 2(qp1)2(q-p-1) because there are 2 choices for $i\notin S$, then qp1q-p-1 choices for $j$, and then $U=S\sqcup\{i\}$. In a type 2 epoch, we have two different contributions to count. First consider $U$ with $|U\cap[p+1]|=p$, and recall that these are never allowed to remove $j$. This gives 2(qp1)2(q-p-1) tuples $(i,j,U)$, since there are again $2$ choices for $i\notin S$, then qp1q-p-1 choices for $j$, and $U=S\sqcup\{i\}$. We also need to consider $U$ with $|U\cap[p+1]|=p-1$ (i.e., cases where $j$ is removed). This gives (p1)(qp2)(p-1)(q-p-2) tuples $(i,j,U)$, since there are p1p-1 choices for $i\in S$, then qp2q-p-2 choices for $j\notin S$, and $U=S\sqcup\{j\}$ (we need $i\in S$ so that $i\in U$, as required by (34)). In total, the number of times that $S$ is used is
@@
Now consider the remaining case: fix $S$ with $h:=|S\cap[p+1]|<p-1$. In a type 1 epoch $\ell$, the number of tuples $(i,j,U)$ assigned to $S$ is (p+1h)(qp1)(p+1-h)(q-p-1) because there are p+1hp+1-h choices for $i\notin S$, then qp1q-p-1 choices for $j$, and $U=S\sqcup\{i\}$. In a type 2 epoch, we again have two contributions to count. First consider $U$ with $|U\cap[p+1]|=h+1$ (i.e., $i$ is removed). This gives (p+1h)(q2p+h1)(p+1-h)(q-2p+h-1) tuples $(i,j,U)$, since there are p+1hp+1-h choices for $i\notin S$, then q2p+h1q-2p+h-1 choices for $j\notin S$, and $U=S\sqcup\{i\}$. Now consider $U$ with $|U\cap[p+1]|=h$ (i.e., $j$ is removed). This gives h(q2p+h1)h(q-2p+h-1) tuples $(i,j,U)$, since there are $h$ choices for $i\in S$, then q2p+h1q-2p+h-1 choices for $j\notin S$, and $U=S\sqcup\{j\}$. In total, the number of times that $S$ is used is
@@ 4.1 Matrix Decomposition Lemma
Let $s$ be the symbolic rank of $A(a)$, which is also the generic rank of $M(a^{(1)}\otimes b^{(1)}\otimes c^{(1)})$. The entries of $A(a)$ are degree-$1$ homogeneous polynomials in $a$, so by Lemma 4.1 there are vectors of polynomials $f_{1}(a),\ldots,f_{s}(a)$ and $g_{1}(a),\ldots,g_{s}(a)$ such that
@@
where we write, e.g., $H_{0}[f_{t}]$ instead of $H_{0}[f_{t}(a)]$ to emphasize that this is a constant (degree-0) vector. Now we can write
@@
The variables $x$ are naturally partitioned into $X_{1}:=\{a_{i}\,:\,i\in[n_{1}]\}$, $X_{2}:=\{b_{i}\,:\,i\in[n_{2}]\}$, and $X_{3}:=\{c_{i}\,:\,i\in[n_{3}]\}$. For $S\subseteq[3]$, let $H_{S}(\cdot)$ denote the terms that are degree $1$ in $X_{i}$ for each $i\in S$ and degree 0 in $X_{i}$ for each $i\notin S$. Note that all terms in $M(a\otimes b\otimes c)$ are degree $1$ in each of $X_{1},X_{2},X_{3}$, so we can rewrite the above as
@@
Now, for any setting of the variables $x^{(\ell)}$, we bound the rank of each $M_{S}$. First, $\mathrm{rank}(M_{\emptyset})\leq s$ because $\mathrm{colspan}(M_{\emptyset})\subseteq\mathrm{span}\{H_{\emptyset}[f_{t}]\,:\,t\in[s]\}$, noting that $H_{\emptyset}[f_{t}(x)]$ is a constant (degree-0) vector, which we denote $H_{\emptyset}[f_{t}]$. Next, note that $H_{\{1\}}[f_{t}(x)]$ has only “aa” variables (not “bb” or “cc”), which we denote by $H_{\{1\}}[f_{t}(x^{(\ell)})]=H_{\{1\}}[f_{t}(a=a^{(\ell)})]$. Each $H_{\{1\}}[f_{t}(a=a^{(\ell)})]$ lies in $\mathrm{span}\{H_{\{1\}}[f_{t}(a=e^{(j)})]\,:\,j\in[n_{1}]\}$ where $e^{(j)}$ are unit basis vectors. This means $\mathrm{rank}(M_{\{1\}})\leq sn_{1}$ because $\mathrm{colspan}(M_{\{1\}})\subseteq\mathrm{span}\{H_{\{1\}}[f_{t}(a=e^{(j)})]\,:\,t\in[s],\,j\in[n_{1}]\}$. Similarly, $\mathrm{rank}(M_{\{2\}})\leq sn_{2}$ and $\mathrm{rank}(M_{\{3\}})\leq sn_{3}$. Applying the same argument to the row span, we also have $\mathrm{rank}(M_{\{1,2,3\}})\leq s$, $\mathrm{rank}(M_{\{2,3\}})\leq sn_{1}$, $\mathrm{rank}(M_{\{1,3\}})\leq sn_{2}$, and $\mathrm{rank}(M_{\{1,2\}})\leq sn_{3}$. Putting it together,
@@
The proof above is similar to that of Theorem 4.4 of [EGOW18], where they give a bound of $r\leq 8n$ for $n\times n\times n$ tensors. We now explain the source of our improvement from $8n$ to $6n+2$. When decomposing $M$ into 8 terms $M_{S}$ for $S\subseteq[3]$, the original proof of [EGOW18] bounds the rank of each term by $sn$. We instead keep track of the specific structure of each term and see that two of them, namely $M_{\emptyset}$ and $M_{\{1,2,3\}}$, admit a better bound of $s$.
@@
The variables $b$ are naturally partitioned into $\bigsqcup_{i\in[k],j\in[d]}B_{ij}$ where $B_{ij}:=\{b_{m}^{(i,j)}\,:\,m\in[n_{i}]\}$. For $S\subseteq[k]\times[d]$, let $H_{S}(\cdot)$ denote the terms that are degree $1$ in $B_{ij}$ for each $(i,j)\in S$ and degree 0 in $B_{ij}$ for each $(i,j)\notin S$. Note that all terms in (35) are degree $1$ in every $B_{ij}$, so we can rewrite the above decomposition as
@@ Appendix A Trivial Flattenings
Since rank (of the $B$ matrix) cannot exceed generic rank (see Section 2.1.2), it suffices to exhibit a single choice of the variables $a^{(i,\ell)}$ such that the $b^{(\ell)}$ vectors are linearly independent. Provided $r\leq\prod_{i\in S}n_{i}$, we can choose the variables $a^{(i,\ell)}$ so that the $b^{(\ell)}$ are distinct standard unit basis vectors. Formally, fix an injection $r]\to\prod_{i\in S}[n_{i}]$ and write $\ell\mapsto(\phi_{i}(\ell)\,:\,i\in S)$ where $\phi_{i}(\ell)\in[n_{i}]$. Then for $i\in S$, set $a^{(i,\ell)}$ equal to the standard unit basis vector $e^{(\phi_{i}(\ell))}\in\mathbb{R}^{n_{i}}$, so that $b^{(\ell)}=\bigotimes_{i\in S}e^{(\phi_{i}(\ell))}$. Applying the same argument to the $c^{(\ell)}$ vectors, we also incur the requirement $r\leq\prod_{i\in[k]\setminus S}n_{i}$. ∎